There are 850 douglas fir and ponderosa pine trees in a section of forest bought by sawz logging co. The company paid a average of $300 for each douglas fir and $225 for each ponderosa pine. if the company paid $217500 for the trees, how many of each kind did the company buy. I wasnt sure at all how to get this one started. i am not sure if which number i am trying break down.
300x+225x-850=217,500. that doesnt make logical sense to me. i know it will 300x and 225x. I am not sure where to go from there.

Answer by Paul(988)

Let D. Trees be x
Let P, trees be y
Now; together they equal 850 trees
so: x+y=850 (solve for y)
y=850-x (subsitution)
D. trees cost $300 and P. trees cost $225
together there cost equal $217500 that includes the amount of tree they buy..
Now the amount are unknown so the equation is.
300x+225y=217500 (subsitue for y)
300x+225(850-x)=217500
300x-225x+191250=217500
75x=26250
x=350
y=850-350
y=500.

Hence, there were 350 D. trees bought, and 500 P. trees bought.
Paul.